OpenStudy (anonymous):

the half life of lead is 3.3 hours. The initial amount of the substance if S (subscript) 0 grams. Express the amount of substance S remaining as a function of time t.

OpenStudy (anonymous):

So, what do we know about radioactive decays?

OpenStudy (anonymous):

the graph is an exponential decay

OpenStudy (anonymous):

Indeed. Do you know the basic formula? Or how to derive it?

OpenStudy (anonymous):

is it -e^?

OpenStudy (anonymous):

Well, it will end up being a function involving e with a negative exponent, but there's more to it.

OpenStudy (anonymous):

I'm not certain what level of math you are at, so I'm just wondering if you were given a formula, or if you are expected to derive it.

OpenStudy (anonymous):

Every 3.3 hours, the amount of S is reduced by half.

OpenStudy (anonymous):

no we were expected to derive it but its from last ur so i cant seem to recall it. haha

OpenStudy (anonymous):

So how much of \(S_0\) is there after 3.3 hours?

OpenStudy (anonymous):

Alright, have you had differential equations?

OpenStudy (anonymous):

It should be \[3.3So/2t\]

OpenStudy (anonymous):

No

OpenStudy (anonymous):

No Not at all.

OpenStudy (anonymous):

Yes i know, but why no, please?

OpenStudy (anonymous):

Every 3.3 hours you are multiplying \(S_0\) by a factor of 1/2.

OpenStudy (anonymous):

Radioactive decay is a random time process where, on average, half of the substance will decay in the half-life time. It's an exponential decay that will reference the amount present, the half-life time, and the time elapsed.

OpenStudy (anonymous):

So \[S = S_0 (\frac{1}{2})^{\frac{t}{3.3}}\]

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

So the rate itself is changing

OpenStudy (anonymous):

Indeed, the rate changes as the material decays. Which is reflected with an exponential decay.

OpenStudy (anonymous):

Oops, polpak, that's what i wanted to write! sorry

OpenStudy (anonymous):

but as vandreigan says, the rate is not constant.

OpenStudy (anonymous):

So Vandreigan: Isn't what polpak wrote is the rate at every instant? Integrating it would do the job, isnt it? Or i'm missing something?

OpenStudy (anonymous):

There are some things you'll need to know that aren't readily apparent, however. A half-life equals Tln2. Find T. \[N(t) = N _{0}e ^{-t/T}\]

OpenStudy (anonymous):

So, since you are using S as the mass, just change N to S, find T, plug it all in. If you need to derive this equation, let me know.

OpenStudy (anonymous):

ok. Thanks All!

OpenStudy (anonymous):

What I wrote was correct.. If you want to rewrite \[(\frac{1}{2})^{\frac{1}{3.3}}\] as a power of e that's fine, but either way it'll be the same number.

OpenStudy (anonymous):

What you wrote has no bearing on the question itself. I'm sorry, but its tangential to the proper approach.

OpenStudy (anonymous):

No, according to you, in 3.3 hours the material should half decay, but its not the case. What i think its the rate at every instant, I'm not sure if its write or wrong.

OpenStudy (anonymous):

sorry right*

OpenStudy (anonymous):

It's an average rate. Decays are rooted in probability.

OpenStudy (anonymous):

Anyway, have a good afternoon. There is plenty of information online involving the solutions to the decay ODE's and the like. Most physics books will cover it pretty well as well!

OpenStudy (anonymous):

Good night, its 1:00 AM here. toodles!

OpenStudy (anonymous):

So \(t_{1/2} = \tau \cdot ln2 \)\[\implies 3.3 = \tau \cdot ln2 \]\[\implies \frac{1}{\tau} = \frac{ln2}{3.3} \]\[ \implies \frac{-1}{\tau} = \frac{-ln2}{3.3}\]\[\implies e^{\frac{-1}{\tau}} = (e^{ln2})^{\frac{-1}{3.3}} = 2^{\frac{-1}{3.3}} = (\frac{1}{2})^{\frac{1}{3.3}}\] Which is exactly what I said. I understand the relationship between exponential decay and ODE's, but this student is coming from an algebraic background and won't really benifit from your explanation as much as just setting up the problem in a straightforward way.